# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ Abstract.
G. R. is funded by the European Commission under the grant agreement n. 101108575 (HORIZON-MSCA project QuantMod).
@@ 1.1. Wild Riemann surfaces
Wild (families of) Riemann surfaces, first introduced by Boalch [Boa14a], are central objects for the study of irregular isomonodromy. Classically, for a pointed compact Riemann surface $(\Sigma,\underline{a})$, where $\underline{a}=(a_{1},\ldots,a_{m})\in\Sigma^{m}$ is an ordered $m$-tuple of distinct points, one can look at the character variety $\mathcal{M}_{Betti}$ parametrising representations of $\pi_{1}(\Sigma\smallsetminus\underline{a})$ valued in a complex reductive algebraic group $G$. When $(\Sigma,\underline{a})$ varies in a holomorphic family over a connected complex manifold $B$, one obtains a family of character varieties over $B$, and an action of the fundamental group of $B$ on $\mathcal{M}_{Betti}$. Fixing the genus $g$ of the Riemann surface and the number $m$ of marked points and considering the resulting moduli stack of pointed Riemann surfaces, one is led to the action of mapping class groups on character varieties, subject of much research (cf. [Wen11], [Pal14], [GPW20], [GLX21] among many others).
@@
Recall that representations $\pi_{1}(\Sigma\smallsetminus\underline{a})\rightarrow G$ correspond to $G$-bundles on $\Sigma$ with a meromorphic connection having regular singularities at $\underline{a}$. In a series of works [Boa01], [Boa02], [Boa07], [Boa12], [Boa14a], Boalch gradually extended the above theory to the more general setting of irregular singularities. The starting point was [Boa01], which studies generic meromorphic connections on trivial vector bundles on the projective line, and shows that the Jimbo–-Miwa–Ueno isomonodromic deformation equations [JMU81] arise from irregular isomonodromy connections, generalising the non-abelian Gauss–Manin connections (cf. the introduction and Theorem 7.1 in [Boa01]).
@@
This provided the first motivation for the general definition of wild Riemann surfaces and wild character varieties, given in [Boa14a]. More precisely, in loc. cit. on the one hand Boalch introduced wild character varieties, generalising the spaces $\mathcal{M}_{Betti}$; on the other hand, he defined admissible families of wild Riemann surfaces over a base $B$, i.e., pointed families of Riemann surfaces over $B$ together with an admissible family of irregular types (whose definition will be recalled below). The main result of [Boa14a] (Theorem 10.2) asserts that, given an admissible family of wild Riemann surfaces over a base $B$, the resulting wild character varieties assemble into a local system of Poisson varieties over $B$, with a complete flat Ehresmann connection (the irregular isomonodromy connection). In particular, one obtains an action of the fundamental group of $B$ on the relevant wild character variety. As envisaged in [Boa18, §3] and [Boa14b, §8] (to which we refer the reader for more background and references), constructing moduli stacks of (admissible families of) wild Riemann surfaces would enable to define wild mapping class groups and study their action on wild character varieties.
@@ 1.1.1. Irregular types
Let $\mathfrak{t}\subset\mathfrak{g}$ be a Cartan sub-algebra of a complex reductive Lie algebra; fix integers $p\geq 1$ and $g\geq 0$. Given a compact Riemann surface $\Sigma$ of genus $g$ and a point $a\in\Sigma$ we consider the completed local ring $\hat{\mathcal{O}}_{\Sigma,a}$ of $\Sigma$ at $a$, with maximal ideal $\hat{\mathfrak{m}}_{\Sigma,a}$. We denote by $\mathscr{T}_{\Sigma,a}^{\leq p}$ the quotient $\hat{\mathfrak{m}}_{\Sigma,a}^{-p}/\hat{\mathcal{O}}_{\Sigma,a}$; it is the $\mathbf{C}$-vector space of germs of meromorphic functions at $a$ with pole order at most $p$, up to homolorphic functions. An (untwisted) irregular type with pole order bounded by $p$ at $a$ is an element $Q\in\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}^{\leq p}_{\Sigma,a}$ [Boa14a, Definition 7.1].
@@ 1.1.2. Families of wild Riemann surfaces
Let $B$ be a complex manifold and $m\geq 1$ an integer. A $B$-family of wild Riemann surfaces of genus $g$ with $m$ marked points, and with pole orders at most $p$, is a triple $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ where $\pi\colon\Sigma\rightarrow B$ is a holomorphic family of compact Riemann surfaces of genus $g$, $\underline{a}=(a_{1},\ldots,a_{m})$ is an $m$-tuple of non-intersecting holomorphic sections of $\pi$, and $\underline{Q}=(Q_{1},\ldots,Q_{m})$ is an $m$-tuple of families of irregular types $Q_{i}(b)\in\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}^{\leq p}_{\pi^{-1}(b),a_{i}(b)}$, smoothly varying with $b\in B$ (cf. [Boa14a, Definition 10.1], [DRT, Definition 1.1] and Definition 2.4.1 below).
@@ 1.2. Aim of the text
For every complex manifold $B$, we consider the groupoid $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B)$ of $B$-families of wild Riemann surfaces with $m$ marked points and pole orders at most $p$ (morphisms being isomorphism commuting with the projection to $B$, respecting the sections and the irregular types). In the body of the text, we will first show that the assignment $B\mapsto\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B)$ is an analytic stack, and the natural map $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}\to\mathsf{M}_{g,m}$ to the stack of $m$-pointed genus $g$ compact Riemann surfaces is (representable by) a vector bundle. Subsequently, we will study the substacks of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$ obtained fixing the order of the poles of the irregular types after evaluation at the roots of $\mathfrak{g}$. The fundamental groups of these substacks will be the (global) wild mapping class groups whose definition is one of the main aims of this text.
@@ 1.3. General notation and terminology
The following notation and conventions will be in force throughout the document, unless otherwise stated.
  • Fix a reductive complex Lie algebra $\mathfrak{g}$ with root system $\Phi$, and a Cartan subalgebra $\mathfrak{t}$. Let $r$ be the dimension of $\mathfrak{t}$ as a complex vector space.
  • Fix integers $p,m\geq 1$ and $g\geq 0$.
  • The category of complex manifolds is denoted by $\mathrm{Man}_{\mathbf{C}}$. For every complex manifold $B$, the tensor product $\mathfrak{t}\otimes_{\mathbf{C}}\Gamma(B,\mathcal{O}_{B})$ is identified naturally with the set of holomorphic functions $f\colon B\rightarrow\mathfrak{t}$. Hence, the contravariant functor from $\mathrm{Man}_{\mathbf{C}}$ to Sets
    $B\mapsto\mathfrak{t}\otimes_{\mathbf{C}}\Gamma(B,\mathcal{O}_{B})$
    is representable by $\mathfrak{t}$.
Further terminology and background on stacks can be found in the appendix.
@@ 2. Wild Riemann surfaces and their moduli
In this section, we will first recall the definition of family of wild Riemann surfaces over a complex manifold $B$; then we will define and study their moduli space. The letter $B$ will always denote a complex manifold.
@@ 2.1. Families of Riemann surfaces and their sections
Let us start by defining families of (compact) Riemann surfaces of genus $g$ over a complex manifold $B$; we will refer to them as Riemann surfaces over $B$.
@@
A Riemann surface over $B$ of genus $g$ is a proper holomorphic submersion $\pi\colon\Sigma\rightarrow B$ of complex manifolds, such that all the fibres are connected Riemann surfaces of genus $g$.
@@ 2.1.2. Local description of sections
If $\sigma\colon B\rightarrow\Sigma$ is a section of a Riemann surface over $B$, then the differential of $\sigma$ is injective at every point $b\in B$. As a consequence of the implicit function theorem, one shows (cf. [Huy05, Corollary 1.1.12]) that, in suitable local charts around $b$ and $\sigma(b)$, the map $\sigma$ has the form $(z_{1},\ldots,z_{n})\mapsto(z_{1},\ldots,z_{n},0)$. In particular $\sigma(B)$ is an effective Cartier divisor: the sheaf of $\mathcal{O}_{\Sigma}$-ideals $\mathcal{I}_{\sigma}$ consisting of functions vanishing on $\sigma(B)$ is locally free of rank one. It sits in a short exact sequence
$0\rightarrow\mathcal{I}_{\sigma}\rightarrow\mathcal{O}_{\Sigma}\rightarrow\sigma_{*}\mathcal{O}_{B}\rightarrow 0$
Furthermore, letting $\mathcal{O}_{\Sigma}(\sigma(B))=\mathcal{H}om_{\mathcal{O}_{\Sigma}}(\mathcal{I}_{\sigma},\mathcal{O}_{\Sigma})$, the evaluation map $\mathcal{I}_{\sigma}\otimes_{\mathcal{O}_{\Sigma}}\mathcal{O}_{\Sigma}(\sigma(B))\rightarrow\mathcal{O}_{\Sigma}$ is an isomorphism.
For an integer $k\geq 1$, let $\mathcal{I}_{\sigma}^{k}$ be $k$-fold product of the ideal sheaf $\mathcal{I}_{\sigma}$; it is isomorphic to the $k$-fold tensor product of $\mathcal{I}_{\sigma}$ (over $\mathcal{O}_{\Sigma}$), because the latter is locally free. The inverse of $\mathcal{I}_{\sigma}^{k}$, denoted by $\mathcal{O}_{\Sigma}(k\sigma(B))$, is the $k$-fold tensor product of $\mathcal{O}_{\Sigma}(\sigma(B))$.
@@ 2.2. Sheaves of tails of meromorphic functions with bounded pole along a section
Fix an integer $k\geq 1$, and a Riemann surface $\pi:\Sigma\rightarrow B$ over $B$ with a section $\sigma$. Boalch introduced the notion of irregular type on $\Sigma$ with a pole at $\sigma$ of order at most $k$, cf. [Boa14a, §10]. The aim of this section is to reformulate the definition in sheaf-theoretic language, which is convenient for the purposes of this article. In particular, this reformulation allows to give a simple proof of Proposition 2.3.2 below, on which our subsequent arguments are based.
@@
Tensoring the exact sequence $0\rightarrow\mathcal{I}_{\sigma}^{k}\rightarrow\mathcal{O}_{\Sigma}\rightarrow\mathcal{O}_{\Sigma}/\mathcal{I}_{\sigma}^{k}\rightarrow 0$ with $\mathcal{O}_{\Sigma}(k\sigma(B))$ which preserves exactness because $\mathcal{O}_{\Sigma}(k\sigma(B))$ is locally free - we obtain a short exact sequence
$0\rightarrow\mathcal{O}_{\Sigma}\rightarrow\mathcal{O}_{\Sigma}(k\sigma(B))\rightarrow\mathscr{T}_{\Sigma,\sigma}^{\leq k}\rightarrow 0$
whose cokernel $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ will be called the sheaf of tails of meromorphic functions on $\Sigma$ with poles along $\sigma(B)$, of order bounded by $k$.
2.2.2.
For example, if $\Sigma$ is a Riemann surface and $\sigma:\{*\}\rightarrow\Sigma$ is the map with image a point $a\in\Sigma$, let $\mathfrak{m}_{\Sigma,a}$ be the maximal ideal of the local ring $\mathcal{O}_{\Sigma,a}$. Then $\mathscr{T}_{\Sigma,a}^{\leq k}$ is the skyscraper sheaf at $a$ with stalk $\mathfrak{m}_{\Sigma,a}^{-k}/\mathcal{O}_{\Sigma,a}\simeq\hat{\mathfrak{m}}_{\Sigma,a}^{-k}/\hat{\mathcal{O}}_{\Sigma,a}$. For a family $\Sigma\rightarrow B$ with a section $\sigma$, choosing charts around a point $b\in B$ and $\sigma(b)\in\Sigma$ as in section 2.1.2, meromorphic functions of the form $\frac{f}{z_{n+1}^{k}}$ with $f$ holomorphic give rise to sections of $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ (in a neighbourhood of $\sigma(b)$). In general, sections of $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ can be written in this form only locally.

### v2

@@ Abstract.
During this work, J. D. has been funded by FCiências.ID, and by the PNRR Grant CF 44/14.11.2022, ‘Cohomological Hall Algebras of Smooth Surfaces and Applications’, led by O. Schiffmann.
G. R. is funded by the European Commission under the HORIZON-MSCA project QuantMod (grant n. 101108575).
M. T. carried out part of this work while being a postdoc at Warwick, funded by the ERC grant n. 101001051.
@@ 1.1. Wild Riemann surfaces
Wild (families of) Riemann surfaces, introduced by Boalch [Boa14a], are central objects for the intrinsic (topological) study of isomonodromic deformations of irregular singular meromorphic connections. Classically, for a pointed compact Riemann surface $(\Sigma,\underline{a})$, where $\underline{a}=(a_{1},\ldots,a_{m})\in\Sigma^{m}$ is an ordered $m$-tuple of distinct points, one can look at the character variety $\mathcal{M}_{Betti}$ parametrising representations of $\pi_{1}(\Sigma\smallsetminus\underline{a})$ valued in a complex reductive algebraic group $G$. When $(\Sigma,\underline{a})$ varies in a holomorphic family over a connected complex manifold $B$, one obtains a family of character varieties over $B$, and an action of the fundamental group of $B$ on $\mathcal{M}_{Betti}$. Fixing the genus $g$ of the Riemann surface and the number $m$ of marked points and considering the resulting moduli stack of pointed Riemann surfaces, one is led to the action of mapping class groups on character varieties, subject of much research (cf. [Gol97], [Wen11], [Pal14], [GPW20], [GLX21] among many others).
@@
Recall that representations $\pi_{1}(\Sigma\smallsetminus\underline{a})\rightarrow G$ correspond to $G$-bundles on $\Sigma$ with a meromorphic connection having regular singularities at $\underline{a}$. In a series of works [Boa01], [Boa02], [Boa07], [Boa12], [Boa14a], Boalch gradually extended the above theory to the more general setting of irregular singularities, building on previous works [BJL79], [MR91], [LR94] (further references and remarks on the history of the subject can be found in the introductions of [Boa21] and [BDR25]). The starting point was [Boa01], which studied generic meromorphic connections on vector bundles on the projective line, showed that the Jimbo–Miwa–Ueno isomonodromic deformation equations [JMU81] arise from flat symplectic Ehresmann connections (irregular isomonodromy connections), and noted that they are generalisations of the nonabelian Gauß–-Manin connections (cf. the introduction and Theorem 7.1 in [Boa01]).
This led to the wild generalisation of nonabelian Hodge theory on curves [Sab99], [BB04] and in turn provided the first motivation for the definition of wild Riemann surfaces and wild character varieties (in the untwisted case, which is our focus here), given in [Boa14a].
@@
More precisely, in loc. cit., on the one hand Boalch constructed (algebraically) general wild character varieties, generalising the spaces $\mathcal{M}_{Betti}$ (cf. [Boa14a, § 8], extending the algebraic construction in the generic setting of [Boa07] and complementing the analytic constructions of [Boa01], [BB04]). On the other hand, he defined admissible families of wild Riemann surfaces over a base $B$, i.e., pointed families of Riemann surfaces over $B$ together with an admissible family of irregular types (whose definition will be recalled below). The main result of [Boa14a] (Theorem 10.2) asserts that, given an admissible family of wild Riemann surfaces over a base $B$, the resulting wild character varieties assemble into a local system of Poisson varieties over $B$, with a complete flat Ehresmann connection - the irregular isomonodromy connection. In particular, one obtains an action of the fundamental group of $B$ on the relevant wild character variety. As envisaged in [Boa18, § 3] and [Boa14b, § 8], constructing moduli stacks of (admissible families of) wild Riemann surfaces would enable to define wild mapping class groups and study their action on wild character varieties.
@@ 1.1.1. Irregular types
Let $\mathfrak{t}\subset\mathfrak{g}$ be a Cartan subalgebra of a finite-dimensional complex reductive Lie algebra; fix integers $p\geq 0$ and $g\geq 0$. Given a compact Riemann surface $\Sigma$ of genus $g$ and a point $a\in\Sigma$ we consider the completed local ring $\hat{\mathcal{O}}_{\Sigma,a}$ of $\Sigma$ at $a$, with maximal ideal $\hat{\mathfrak{m}}_{\Sigma,a}$. We denote by $\mathscr{T}_{\Sigma,a}^{\leq p}$ the quotient $\hat{\mathfrak{m}}_{\Sigma,a}^{-p}/\hat{\mathcal{O}}_{\Sigma,a}$: it is the $\mathbf{C}$-vector space of germs of meromorphic functions at $a$ with pole order at most $p$, up to holomorphic terms. An (untwisted) irregular type with pole order bounded by $p$ at $a$ is an element $Q\in\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}^{\leq p}_{\Sigma,a}$ [Boa14a, Definition 7.1].
@@ 1.1.2. Families of wild Riemann surfaces
Let $B$ be a complex manifold, $m\geq 1$ an integer, and $\underline{p}=(p_{1},\ldots,p_{m})\in\mathbf{Z}_{\geq 0}^{m}$ an $m$-tuple of nonnegative integers. A $B$-family of wild Riemann surfaces of genus $g$ with $m$ marked points, and with pole orders at most $\underline{p}$, is a triple $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ where $\pi\colon\Sigma\rightarrow B$ is a holomorphic family of compact Riemann surfaces of genus $g$, $\underline{a}=(a_{1},\ldots,a_{m})$ is an $m$-tuple of mutually non-intersecting holomorphic sections of $\pi$, and $\underline{Q}=(Q_{1},\ldots,Q_{m})$ is an $m$-tuple of families of irregular types $Q_{i}(b)\in\mathfrak{t}\otimes_{\mathbf{C}}\mathscr{T}^{\leq p_{i}}_{\pi^{-1}(b),a_{i}(b)}$, holomorphically varying with $b\in B$ (cf. [Boa14a, Definition 10.1], [DRT, Definition 1.1] and Definition 2.4.1 below).
@@ 1.2. Aim of the text
For every complex manifold $B$, we consider the groupoid $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B)$ of $B$-families of wild Riemann surfaces with $m$ marked points and pole orders at most $\underline{p}$ (morphisms being isomorphisms commuting with the projection to $B$, respecting the sections and the irregular types). In the body of the text, we will first show that the assignment $B\mapsto\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B)$ is an analytic stack, and the natural map $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\longrightarrow\mathsf{M}_{g,m}$ to the stack of $m$-pointed genus $g$ compact Riemann surfaces is (representable by) a vector bundle. Subsequently, we will study the substacks $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ obtained fixing the order $\underline{\mathbf{d}}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{m})$ of the poles of the irregular types after evaluation at the roots of $\mathfrak{g}$. The fundamental groups of these substacks will be the (global) wild mapping class groups whose definition is one of the main aims of this text. Finally, in the last part of the document we will describe more explicitly $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ and $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$ in some examples; we will then define algebraic versions of these stacks, and show that they are Deligne–Mumford if and only if $2g-2+m+\sum_{i}\max(\mathbf{d}_{i})>0$ - as foreseen in [BDR25, p. 399] - generalising the analogous result for moduli stacks of curves.
@@ 1.3. General notation and terminology
The following notation and conventions will be in force throughout the document, unless otherwise stated:
  • fix a finite-dimensional complex reductive Lie algebra $\mathfrak{g}$ and a Cartan subalgebra $\mathfrak{t}$, and let $\Phi$ be the associated root system;
  • let $r$ be the dimension of $\mathfrak{t}$ as a complex vector space;
  • fix integers $m\geq 1$ and $g\geq 0$, and a tuple $\underline{p}=(p_{1},\ldots,p_{m})\in\mathbf{Z}_{\geq 0}^{m}$;
  • the category of complex manifolds (and holomorphic maps) is denoted by $\mathrm{Man}_{\mathbf{C}}$; the letter $B$ will always denote a complex manifold.
Further terminology and background on stacks can be found in appendix A.
@@ 2. Wild Riemann surfaces and their moduli
In this section, we will first recall the definition of families of wild Riemann surfaces over a complex manifold $B$. Then we will define and study their moduli space.
@@ 2.1. Families of Riemann surfaces and their sections
Let us start by defining families of (compact) Riemann surfaces of genus $g$ over a complex manifold $B$. We will call them Riemann surfaces over $B$.
@@
A Riemann surface over $B$ of genus $g$ is a proper holomorphic submersion $\pi\colon\Sigma\rightarrow B$ of complex manifolds, such that all the fibres are Riemann surfaces of genus $g$.
@@ 2.1.2. Local description of sections
If $\sigma\colon B\rightarrow\Sigma$ is a (holomorphic) section of a Riemann surface over $B$, then the differential of $\sigma$ is injective at every point $b\in B$. As a consequence of the implicit function theorem, one shows (cf. [Huy05, Corollary 1.1.12]) that, in suitable local charts around $b$ and $\sigma(b)$, the map $\sigma$ has the form $(z_{1},\ldots,z_{n})\mapsto(z_{1},\ldots,z_{n},0)$. In particular $\sigma(B)$ is an effective Cartier divisor: the sheaf of $\mathcal{O}_{\Sigma}$-ideals $\mathcal{I}_{\sigma}$ consisting of functions vanishing on $\sigma(B)$ is locally free of rank one. It sits in a short exact sequence
$0\rightarrow\mathcal{I}_{\sigma}\rightarrow\mathcal{O}_{\Sigma}\rightarrow\sigma_{*}\mathcal{O}_{B}\rightarrow 0.$
Furthermore, letting $\mathcal{O}_{\Sigma}(\sigma)=\mathcal{H}om_{\mathcal{O}_{\Sigma}}(\mathcal{I}_{\sigma},\mathcal{O}_{\Sigma})$, the evaluation map $\mathcal{I}_{\sigma}\otimes_{\mathcal{O}_{\Sigma}}\mathcal{O}_{\Sigma}(\sigma)\rightarrow\mathcal{O}_{\Sigma}$ is an isomorphism.
For an integer $k\geq 0$, let $\mathcal{I}_{\sigma}^{k}$ be the $k$-fold product of the ideal sheaf $\mathcal{I}_{\sigma}$; it is isomorphic to the $k$-fold tensor product of $\mathcal{I}_{\sigma}$ (over $\mathcal{O}_{\Sigma}$), because the latter is locally free. The inverse of $\mathcal{I}_{\sigma}^{k}$, denoted by $\mathcal{O}_{\Sigma}(k\sigma)$, is isomorphic to the $k$-fold tensor product of $\mathcal{O}_{\Sigma}(\sigma)$.
@@ 2.2. Sheaves of tails of meromorphic functions with bounded pole along a section
Fix an integer $k\geq 0$, and a Riemann surface $\pi\colon\Sigma\rightarrow B$ over $B$ with a section $\sigma$. In [Boa14a, § 10] Boalch considers irregular types on $\pi^{-1}(b)$ with a pole at $\sigma(b)$, varying smoothly with $b\in B$.
The aim of this section is to express in sheaf-theoretic terms the notion of “holomorphic family of irregular types” on $\Sigma$ with a pole at $\sigma$ of order at most $k$. This will allow us to give a simple proof of Proposition 2.3.4 below, on which our subsequent arguments are based.
@@
Tensoring the exact sequence $0\rightarrow\mathcal{I}_{\sigma}^{k}\rightarrow\mathcal{O}_{\Sigma}\rightarrow\mathcal{O}_{\Sigma}/\mathcal{I}_{\sigma}^{k}\rightarrow 0$ with $\mathcal{O}_{\Sigma}(k\sigma)$ we obtain a short exact sequence
$0\rightarrow\mathcal{O}_{\Sigma}\rightarrow\mathcal{O}_{\Sigma}(k\sigma)\rightarrow\mathscr{T}_{\Sigma,\sigma}^{\leq k}\rightarrow 0,$
whose cokernel $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ will be called the sheaf of tails of meromorphic functions on $\Sigma$ with pole along $\sigma(B)$, of order bounded by $k$. This object is central to our construction: on the one hand, we will use it to define (untwisted) irregular types in section 2.3; on the other hand, exponential factors of such irregular types are global sections of $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$, cf. section 2.4.3.
2.2.2 Example.
  1. (1)
    Let $\Sigma$ be a compact Riemann surface and let $\sigma\colon\{*\}\rightarrow\Sigma$ be the map with image a point $a\in\Sigma$. Let $\mathfrak{m}_{\Sigma,a}$ be the maximal ideal of the local ring $\mathcal{O}_{\Sigma,a}$. Then $\mathscr{T}_{\Sigma,a}^{\leq k}$ is the skyscraper sheaf at $a$ with stalk $\mathfrak{m}_{\Sigma,a}^{-k}/\mathcal{O}_{\Sigma,a}\simeq\hat{\mathfrak{m}}_{\Sigma,a}^{-k}/\hat{\mathcal{O}}_{\Sigma,a}$.
  2. (2)
    For a family $\Sigma\rightarrow B$ with a section $\sigma$, choosing charts around a point $b\in B$ and $\sigma(b)\in\Sigma$ as in section 2.1.2, meromorphic functions of the form $\frac{f}{z_{n+1}^{k}}$ with $f\colon\mathbf{C}^{n+1}\rightarrow\mathbf{C}$ holomorphic give rise to sections of $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ in a neighbourhood of $\sigma(b)$. In general, sections of $\mathscr{T}_{\Sigma,\sigma}^{\leq k}$ can be written in this form only locally.